OpenStudy (anonymous):

the population of the world in billions of people can be modeled by the function f(t) = 5.3(1.018)^t where t is years since 1990 find f(0) and f'(0) find f(30) and f'(30) using units explain what each answer tells you about the population of the world

OpenStudy (anonymous):

f(0) = 5.3 billions f'(t) = 5.3 ln (1.018) (1.018) ^t = .095 * (1.018) ^t => f'(0) = .095 billions f(30) = 5.3 ( 1.018) ^30 = 9.05 billions f'(30) = 5.3 ln (1.018) (1.018) ^30 = .16 billions The population keeps growing with the rate of .16 billion for 30 years.

OpenStudy (anonymous):

THANK YOU!!!!

OpenStudy (anonymous):

Do you have any question?

OpenStudy (anonymous):

I think that covered it... I had to think about it for a minute but I got it... I have a guestion about something else if you don't mind

OpenStudy (anonymous):

Estimate f'(2) for f(x) =3^x ... explain reasoning